OpenStudy (anonymous):

What is the value of 54 – 4a^2 + 3b^3 when a = −2 and b = 4?

OpenStudy (anonymous):

Just plug the values in. Where are you have troubles?

OpenStudy (anonymous):

\[54 - 4(-2)^2 + 3(4^3)\]

OpenStudy (anonymous):

*having

OpenStudy (anonymous):

\[54-4(-2)^2+3(4^3) = 54-4*4+3*64\]

OpenStudy (anonymous):

Yeah, exactly.

OpenStudy (anonymous):

Is 106 correct

OpenStudy (anonymous):

Er, what?

OpenStudy (anonymous):

For the final answer

OpenStudy (anonymous):

4*4 = 16, 3*63 = 192

OpenStudy (anonymous):

52-16+192 = 230

OpenStudy (anonymous):

*54

OpenStudy (anonymous):

230 then

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ty